OpenStudy (anonymous):

write the equation of the line that has a slope of -(3/4) and passes through (-2,1)

OpenStudy (anonymous):

please solve

OpenStudy (anonymous):

y=-3/4x-1/2 i plugged in -2 for x -3/4(-2)-1/2=1

OpenStudy (ybarrap):

y = mx + b 1 = (-3/4) * (-2) + b b = 1 - 3/2 = -1/2 y = (-3/4)x - 1/2 Ans. -3x/4 - 1/2